If the APR for a car loan is 5%, what is the effective annual interest rate (in %) if interest on the loan is compounded monthly and you have biweekly payments?

Is the answer 0.20812%?
Answer #1

EFFECTIVE RATE OF INTEREST IF INTEREST IS PAID ON MONTHLY BASIS:

=(1+i/n)n-1

=(1+0.05/12)12-1

=1.051162-1

=0.051162

=5.1162%

EFFECTIVE RATE OF INTEREST IF INTEREST IS PAID ON WEEKELY BASIS:

=(1+i/n)n-1

=(1+0.05/52)52-1

=1.051246-1

=0.051246

5.1246%